business finance

What is the present value of

$

to be received

years from now discounted back to the present at

percent?

The present value of

$

to be received

years from now discounted back to the present at

percent is

$nothing.

(Round to the nearest cent.)
Answer #1

a.Information provided:

Future value= $800

Time= 9 years

Discount rate= 10%

Enter the below in a financial calculator to compute the present value:

FV= 800

N= 9

I/Y= 10

Press the CPT key and PV to calculate the present value.

The value obtained is 339.28.

Therefore, the present value is $339.28.

b.Information provided:

Future value= $400

Time= 8 years

Discount rate= 9%

Enter the below in a financial calculator to compute the present value:

FV= 400

N= 8

I/Y= 9

Press the CPT key and PV to calculate the present value.

The value obtained is 200.75.

Therefore, the present value is $200.75.

c.Information provided:

Future value= $1,100

Time= 11 years

Discount rate= 4%

Enter the below in a financial calculator to compute the present value:

FV= 1,100

N= 11

I/Y= 4

Press the CPT key and PV to calculate the present value.

The value obtained is 714.54.

Therefore, the present value is $714.54.

d.Information provided:

Future value= $1,050

Time= 6 years

Discount rate= 19%

Enter the below in a financial calculator to compute the present value:

FV= 1,050

N= 6

I/Y= 19

Press the CPT key and PV to calculate the present value.

The value obtained is 369.75.

Therefore, the present value is $369.75.


answered by: ANURANJAN SARSAM
Answer #2

SOLUTION :


PV 

= FV / (1 + r)^n

= 500 / ( 1 + 0.08)^8

= 270.13 ($) 


So, PV of $500 to be received in 8 years from now and discounted back to the present at the rate of 8% is = 270.13 ($) (ANSWER)


answered by: Tulsiram Garg